# Universal Approximation Theory: Foundations for Parallelism in Neural Networks

### v1

@@ Abstract
Neural networks are increasingly evolving towards training large models with big data, a method that has demonstrated superior performance across many tasks. However, this approach introduces an urgent problem: current deep learning models are predominantly serial, meaning that as the number of network layers increases, so do the training and inference times. This is unacceptable if deep learning is to continue advancing. Therefore, this paper proposes a deep learning parallelization strategy based on the Universal Approximation Theorem (UAT). From this foundation, we designed a parallel network called Para-Former to test our theory. Unlike traditional serial models, the inference time of Para-Former does not increase with the number of layers, significantly accelerating the inference speed of multi-layer networks. Experimental results validate the effectiveness of this network.
@@ 1 Introduction
In today’s era, deep learning models have undoubtedly taken the lead in the fields of computer vision (CV) and natural language processing (NLP), demonstrating exceptional performance across numerous tasks. In CV, models such as ResNet(He et al. 2015) and ViT(Vaswani et al. 2017) for image recognition, U-Net (Ronneberger, Fischer, and Brox 2015) for image segmentation, GAN (Goodfellow et al. 2014) for image generation, and diffusion models like SORA (Liu et al. 2024) for video generation highlight the extensive application of deep learning in CV. These models are primarily based on convolutional neural networks (CNNs), Transformers, or a combination of both. Simultaneously, large language models (LLMs) based on Transformers have become dominant in natural language processing. Overall, deep learning models have shown vast application potential and powerful capabilities so far.
However, this progress has brought forth some challenges: the development trend in deep learning is moving towards larger models and larger datasets. This trend is fundamentally driven by the UAT(Cybenko 2007; Hornik, Stinchcombe, and White 1989). From the papers UAT2LLM(Wang and Li 2024) and UAT2CV(Wang and Li 2024), we know that residual-based CNNs and Transformer-based models are specific implementations of UAT, both exhibiting the characteristic of dynamically approximating functions based on inputs. Larger models and datasets are indeed requirements of the UAT theory: deeper network layers can more precisely approximate target functions dynamically based on inputs, while larger and more diverse datasets help ensure the training data closely represents real-world conditions, making the model’s approximated functions more reflective of real-world complexities.
@@
However, this trend inevitably leads to increased demands for computational resources and extended training times. As more powerful models are needed, the UAT theory mathematically necessitates larger networks, which in turn require more computational resources. While technological advancements, especially in hardware performance, provide potential ways to alleviate this issue, the increase in time costs remains a fundamental obstacle to scaling serial networks. The improvement in computational resources such as GPUs and TPUs is limited by the speed of electrical signal transmission, whereas the theoretical scale of networks has no such upper limit. When networks reach tens or even hundreds of thousands of layers, inference times can extend to several minutes or even hours for each batch. Current solutions include model optimization techniques like quantization and pruning (Sun et al. 2023; Ma, Fang, and Wang 2023; Xia et al. 2023), hardware acceleration with specialized accelerators and edge computing (Gale et al. 2020), and parallel and distributed computing (Ben-Nun and Hoefler 2018) with model parallelization, data parallelization, and distributed computing frameworks. However, these do not fundamentally address the inference delay caused by the increasing number of serial network layers.
@@
We believe the root cause lies in the serial computation mechanism of existing deep learning models, where the output of one layer serves as the input for the next. This design stems from the early development of deep learning networks in the field of CV, where image data exhibit local correlations and spatiotemporal invariance. To accommodate these features, early CV networks primarily employed convolutional layers with small receptive fields. To achieve a larger receptive field, deeper network layers were necessary, leading to early deep-learning architectures like AlexNet and VGG. The introduction of ResNet brought a paradigm shift, as residual networks have the ability to dynamically fit functions based on input, influencing almost all subsequent network designs and solidifying the tradition of multi-layer stacking.
While this design was manageable with fewer layers, the increasing depth and number of parameters in modern networks have created a critical issue: slow computation efficiency due to the need for sequential layer-by-layer calculations during inference. This computation method inevitably leads to delays, which become more severe as the network depth increases, significantly constraining the progress of CV and NLP technologies.
@@
Therefore, it is imperative to explore parallelization techniques for deep learning networks, enabling parallel computation across multiple layers. This approach is key to accelerating the inference process, breaking the time bottleneck, and advancing CV and NLP technologies towards greater efficiency and speed. To develop such parallel computation, we must start from the fundamental theory of deep learning: the UAT. The network must be designed in accordance with UAT principles and must satisfy the requirements outlined in the papers UAT2LLM and UAT2CV, which emphasize the ability to dynamically fit functions based on data. With this in mind and leveraging existing network structures, we designed a parallel network, Para-Former, to validate the feasibility of our theory. Our contributions include:
@@
The structure of this paper is as follows: In Section 2, we first present the general form of the Universal Approximation Theorem (UAT). Based on the mathematical form of UAT, we define the parallel network Para-Former in Section 3. In Section 4, we validate the effectiveness of Para-Former in various ways: we assess its feasibility in Section 4.2, explore the impact of network depth in Section 4.3, and analyze the influence of different datasets in Section 4.4.
@@ 2 UAT
Deep learning networks have demonstrated their powerful capabilities in numerous CV and NLP tasks. In the papers UAT2LLMs and UAT2CVs, it has been proven that both Transformer-based networks and residual-based CNNs are specific implementations of the UAT. Therefore, we consider UAT to be the fundamental theory for deep learning. To design parallelized networks, we must start from the principles of UAT. Before formally designing parallel networks, we will first review the basic form of the UAT theory and then propose a network parallelization solution based on this mathematical theory. The UAT was initially proposed by Cybenko (Cybenko 2007), and its basic form is as follows:
@@
This indicates that when $N$ is sufficiently large, a neural network can approximate any continuous function on a closed interval. Hornik, Stinchcombe, and White  (Hornik, Stinchcombe, and White 1989) further proved that this function can approximate any Borel measurable function. Observing Equation (1), where the function $G(\mathbf{x})$ produces scalar outputs in the real number field $\mathbb{R}$, the scenario naturally extends when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring approximation in each dimension. To accommodate such multi-dimensional outputs, we simply need to extend Equation (1): the transformation matrix $\mathbf{W}_{j}$ is revised to the space $\mathbb{R}^{n\times m}$, the bias term $\theta_{j}$ is recast as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a matrix. We call the parameters $\mathbf{W}_{j}$ and $\alpha_{j}$ in UAT as weight and $\theta_{j}$ as bias.
@@
3 UAT for Parallel Network
3.1 Pre-request for Basic Block of Parallel Network
From Equation (1), we can observe a significant characteristic: if we define different indices $j$ as the indices of layers and the corresponding layer’s mathematical form is $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$, it is evident that the parameters between different layers do not interact with each other. Therefore, based on (1), we can design networks where parameters across multiple layers are independent, achieving parallel computation, like Figure 1. Figure 1.a represents a typical sequential network, which clearly requires computation to proceed in order from front to back. In contrast, Figure 1.b depicts a parallel network, where there is no parameter interaction between different blocks, allowing them to be considered independent and thus suitable for parallel computation. The blocks in this network should satisfy the formula given in Figure 1.b. Clearly, under this condition, the sum of the multiple blocks’ outputs adheres to the UAT theory.
If we design the parallel network solely based on Eq. (1), the network will face a limitation once training is complete: the function that can be approximated by the network is fixed, as the parameters in the corresponding UAT are fixed. So to ensure the network can dynamically approximate functions based on input, some or all of the parameters $\alpha_{j}$, $W_{j}$ and $\theta_{j}$ must be influenced by the input (referencing to UAT2LLM and UAT2CV), which can be simply written as $\alpha_{j}(\mathbf{x})$, $W_{j}(\mathbf{x})$ and $\theta_{j}(\mathbf{x})$.
This allows the network to dynamically adapt and fit the function based on the input while enabling parallel computation. So we could design the basic block in Figure 1, like $\text{Block}_{1}\cdots\text{Block}_{n}$ are the implementation of $\alpha_{j}(\mathbf{x}))\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}(\mathbf{x})\mathbf{x}+\theta_{j}(\mathbf{x})\right),i=1\cdots n$. Some or all the parameters $\alpha_{j}$, $W_{j}$ and $\theta_{j}$ will change according to input.
@@ 3.1 Pre-request for Basic Block of Parallel Network
Figure 3: The structure of the Para-Former network. a.Para-Former-1-n and b.Para-Former-m-n represent Para-Former with 1 depth of n layers and m depth of n layers, respectively.
3.2 Para-former
$=$ $\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}$
$\hat{\mathbf{H}}_{i}$ $=\operatorname{Attention}(\mathbf{x}_{i}\mathbf{W}_{Qi}=\mathbf{Q}_{i},\mathbf{x}_{i}\mathbf{W}_{Ki}=\mathbf{K}_{i},$ (4)
$\mathbf{x}_{i}\mathbf{W}_{Vi}=\mathbf{V}_{i})$
$=\operatorname{softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{T}}{\sqrt{M}}\right)\mathbf{V}_{i}$
$=\mathbf{H}_{i}\mathbf{V}_{i}=\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$
The above representation mainly describes the engineering process rather than its true mathematical form. According to UAT2LLM, the Multi-Head Attention (MHA) and Feed-Forward Network (FFN) components of Transformers can be expressed in matrix-vector form. We denote the matrix-vector form of variables with a prime symbol (e.g., $\mathbf{x}_{i}$ becomes $\mathbf{x}^{\prime}_{i}$). For more details, refer to UAT2LLM.
$MHA(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}$ (6)
$FFN(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}\$ (7)
The parameter $\mathbf{W}^{\prime}_{i,1}$ is influenced by the input. As shown in Figure 2, the output of a single Transformer layer, $\mathbf{x}^{\prime}_{i+1}$, can be described by Eq. (8) (Details could be found in Section A):
$\mathbf{x}^{\prime}_{i+1}$ $=\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}$ (8)
The mathematical form in Eq. (8) meets the requirements of a basic module in UAT. Therefore, we can use the structure in Figure 2 as the basic module in Para-Former, corresponding to Figure 3.a. Besides, we add a symbol of a hat on some parameters to represent the parameters will be affected by the input, like $\hat{\mathbf{W}^{\prime}_{i,1}}$.
$\mathbf{x}^{\prime}_{i+2}$ $=\hat{\mathbf{W}^{\prime}_{i+1,2}}\mathbf{x}^{\prime}_{i}+\hat{\mathbf{W}^{\prime}_{i+1,3}}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})$ (9)
$+\mathbf{W}^{\prime}_{i+1,3}\sigma\{\hat{\mathbf{W}^{\prime}_{i+1,4}}\mathbf{x}^{\prime}_{i}+\hat{\mathbf{b}^{\prime}_{i+1,4}}\}+\hat{\mathbf{b}^{\prime}_{i+1,5}}$
The output of a two-layer Transformer is shown in Eq. (9) (Details could be found in Section A). From Eq. (9), it is evident that a serial configuration of multi-layer Transformers also conforms to the mathematical form of UAT. Therefore, we can extend the Para-Former structure. That is, within one block in Figure 1, there can be multiple serially connected Transformer modules, while multiple blocks operate in parallel, as shown in Figure 3.b. We refer to the number of Transformer modules within a block as the depth and the total number of Transformer modules across all blocks as the network layer count.
3.3 Speed-up Ratio
We have proposed a network parallelization scheme, and in this section, we will compare the running speed of this parallelized network with that of traditional serial networks. It is important to note that due to differences in hardware and implementation methods, providing an objective comparison through experiments can be challenging. Therefore, we will focus on a theoretical analysis. Based on the network structure shown in Figure 2, the inference time of Para-Former is only affected by its network depth.
Assume we use a Para-Former with a depth of $M$ and a layer count of $L$, compared to a traditional Transformer network with a depth of $N$ layers, where $L>>N$. In this case, the running speed of the parallel network would be $N/M$ times faster than that of the serial network. This is because the speed of the parallel network is solely related to the depth of Para-Former. As the network depth increases, the acceleration effect becomes more pronounced.
4 Experiments
Our experimental approach is to first verify the feasibility of the parallel network theory and its adherence to the UAT approximation properties using several commonly used public image classification datasets (Section 4.1). We then explore the impact of network depth on the UAT theory (Section 4.3), followed by an investigation into the influence of the dataset (Section 4.4).
Dataset Cifar-10 Cifar-100 STL-10 OxfordIIITPet Flower-102
@@ 4 Experiments
4.1 Dataset
In pursuit of a comprehensive understanding of the efficacy and practicality of parallel models, we meticulously selected several widely-acknowledged and representative image classification datasets for an integrated comparative analysis. These datasets encompass a diverse array of visual content and span various levels of difficulty, thereby furnishing us with a fertile ground to probe the performance boundaries of parallel models. Specifically, our study centers on the following pivotal datasets:
1. CIFAR-10 & CIFAR-100 (Krizhevsky 2009): These were constructed by Alex Krizhevsky et al. CIFAR-10 encompasses 60,000 32x32 pixel color images across 10 classes, while CIFAR-100 extends this to 100 categories.
2. STL-10 (Coates, Ng, and Lee 2011): STL-10 comprises 5,000 training and 8,000 testing images distributed among 10 classes, each at a 96x96 pixel resolution.
3. Flowers-102 (Nilsback and Zisserman 2008): Focused on floral categorization, this dataset houses 102 distinct flower species, totalling 8,189 images with a per-species count ranging from 40 to 258.
4. Oxford-IIIT Pets (Parkhi et al. 2012): Tailored for the recognition of pet breeds (cats and dogs), this dataset contains over 7,300 images across 37 categories, characterized by varied animal poses, backgrounds, and lighting conditions.
In Table 1, we provide more details for these dataset.
Through experimentation on these meticulously chosen datasets, we aspire to thoroughly validate the feasibility and strengths of parallel models across a spectrum of scales, complexities, and domains in image classification tasks. Simultaneously, we seek to elucidate potential limitations and chart avenues for future research endeavors.
Table 2: This table shows the predictions of Para-former-1-1, Para-former-1-6, Para-former-1-12, Para-former-1-24 on the datasets CIFAR-10, CIFAR-100, STL-10, OxfordIIITPet, and Flower-102. We simply represent them as Para-F-1-1, Para-F-1-6, Para-F-1-12 and Para-F-1-24, simply.
Dataset Para-F-1-1 Para-F-1-6 Para-F-1-12 Para-F-1-24

### v4

@@ Abstract
Neural networks are increasingly evolving towards training large models with big data, a method that has demonstrated superior performance across many tasks. However, this approach introduces an urgent problem: current deep learning models are serial computing, meaning that as the number of network layers increases, so do the training and inference times. This is unacceptable if deep learning models go deeper. Therefore, this paper proposes a deep learning parallelization strategy based on the Universal Approximation Theorem (UAT). From this foundation, we designed a parallel network called Para-Former to test our theory. Unlike traditional serial models, the inference time of Para-Former does not increase with the number of layers, significantly accelerating the inference speed of multi-layer networks. Experimental results validate the effectiveness of this network.
@@ 1 Introduction
Currently, deep learning is advancing towards larger models and datasets, driven by the success of ChatGPT (Radford and Narasimhan 2018; Brown et al. 2020; Achiam et al. 2023), Llama (Touvron et al. 2023), and ViT (Dosovitskiy et al. 2020), among others. These large models demonstrate significantly enhanced performance compared to their predecessors. According to UAT2LLMs (Wang and Li 2024b) and UAT2CVs (Wang and Li 2024b), residual-based convolutional neural networks (CNNs) and Transformer-based large language models are specific implementations of the UAT (Cybenko 2007; Hornik, Stinchcombe, and White 1989), capable of dynamically approximating functions based on inputs. From the perspective of UAT, larger models inherently possess stronger dynamic fitting capabilities (Wang and Li 2024b, a). Therefore, the trend towards larger models is undoubtedly the future direction of deep learning, and large datasets are essential for training these models to ensure their approximations accurately reflect real-world complexities.
@@
However, this trend towards larger models and datasets inevitably leads to increased demands for computational resources and extended training times. Despite advancements in hardware technology, such as GPUs and TPUs, the rise in time costs remains a significant obstacle. The speed of electrical signal transmission is limited, whereas the theoretical scale of networks has no such upper limit. As network layers increase to tens or even hundreds of thousands, inference times can extend significantly, potentially requiring several minutes or hours for a single batch, making training large models on large datasets exceedingly difficult.
Current solutions include model optimization techniques such as quantization (Jacob et al. 2017; Lu, Zhong, and Yang 2022; Guo, Wang, and Cui 2021; Lin et al. 2023), pruning (Ma, Fang, and Wang 2023; Xia et al. 2023), and knowledge distillation (Gou et al. 2020; Gu et al. 2021). Additionally, specialized accelerators (Jouppi et al. 2017; Elbtity, Chandarana, and Zand 2024) and edge computing (Shi et al. 2016; Gale et al. 2020), as well as parallel and distributed computing (Shallue et al. 2018; Ben-Nun and Hoefler 2018; Sergeev and Balso 2018), offer potential solutions. However, these methods do not fundamentally address the inference delay caused by the increasing number of serial network layers.
@@
The root cause lies in the serial computation mechanism of existing deep learning models, where the output of one layer serves as the input for the next. This design originated from early computer vision networks, which used convolutional layers with small receptive fields to accommodate the local correlations and spatiotemporal invariance of image data. Achieving a larger receptive field required deeper network layers, as exemplified by architectures like AlexNet (Krizhevsky, Sutskever, and Hinton 2012) and VGG (Simonyan and Zisserman 2014). The introduction of ResNe t (He et al. 2015) marked a paradigm shift, as the residual structure could dynamically fit functions based on input (Wang and Li 2024b, a), influencing subsequent network designs and establishing the tradition of multi-layer stacking. While this design was manageable with fewer layers, the increasing depth and number of parameters in modern networks have created a critical issue: slow computation efficiency due to the need for sequential layer-by-layer calculations during inference. This computation method inevitably leads to delays, which become more severe as network depth increases, significantly constraining the progress of computer vision and natural language processing technologies.
@@
Therefore, it is imperative to explore parallelization techniques for deep learning networks, enabling parallel computation across multiple layers. This approach is key to accelerating the inference process, breaking the time bottleneck, and advancing computer vision and natural language processing technologies towards greater efficiency and speed. To develop such parallel computation, we must start from the fundamental theory of deep learning: the UAT. The network must be designed in accordance with UAT principles and must satisfy the requirements outlined in the papers UAT2LLMs (Wang and Li 2024b) and UAT2CVs (Wang and Li 2024a), which emphasize the ability to dynamically fit functions based on data. With this in mind and leveraging existing network structures, we designed a parallel network, Para-Former, to validate the feasibility of our theory. Our contributions include:
@@
The structure of this paper is as follows: In Section 2, we first present the general form of the UAT. Based on the mathematical form of UAT, we define the parallel network Para-Former in Section 3. In Section 4, we validate the effectiveness of Para-Former in various ways: we assess its feasibility in Section 4.2, explore the impact of network depth in Section 4.3, and analyze the influence of datasets on network performances in Section 4.4.
@@ 2 UAT
Deep learning networks have demonstrated their powerful capabilities in numerous CV and NLP tasks. In the papers UAT2LLMs (Wang and Li 2024b) and UAT2CVs (Wang and Li 2024a), they have been proven that both Transformer-based networks and residual-based CNNs are specific implementations of the UAT. Therefore, we consider UAT to be the fundamental theory for deep learning. To design parallelized networks, we must start from the principles of UAT. Before formally designing parallel networks, we will first review the basic form of the UAT theory and then propose a network parallelization solution based on this mathematical theory. The UAT was initially proposed by Cybenko (Cybenko 2007), and its basic form is as follows:
@@
This indicates that when $N$ is sufficiently large, a neural network can approximate any continuous function on a closed interval. Hornik, Stinchcombe, and White  (Hornik, Stinchcombe, and White 1989) further proved that this function can approximate any Borel measurable function. Observing Eq. (1), where the function $G(\mathbf{x})$ produces scalar outputs in the real number field $\mathbb{R}$, the scenario naturally extends when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring approximation in each dimension. To accommodate such multi-dimensional outputs, we simply need to extend Eq. (1): the transformation matrix $\mathbf{W}_{j}$ is revised to the space $\mathbb{R}^{n\times m}$, the bias term $\theta_{j}$ is recast as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a matrix. We call the parameters $\mathbf{W}_{j}$ and $\alpha_{j}$ in UAT as weight and $\theta_{j}$ as bias.
3 UAT for Parallel Network
3.1 Requirements for Basic Block in Parallel Network
From Eq. (1), we can observe a significant characteristic: if we define different indices $j$ as the indices of layers and the corresponding layer’s mathematical form is $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$, it is evident that the parameters between different layers do not interact with each other. Therefore, based on Eq. (1), we can design networks where parameters across multiple layers are independent, achieving parallel computation, like Figure 1. Figure 1.a represents a typical sequential network, which clearly requires computation to proceed in order from front to back. In contrast, Figure 1.b depicts a parallel network, where there is no parameter interaction between different blocks, allowing them to be considered independent and thus suitable for parallel computation. The blocks in this network should satisfy the formula of $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$. Clearly, under this condition, the sum of the multiple blocks’ outputs adheres to the UAT theory.
If we design the parallel network solely based on Eq. (1), the network will face a limitation once training is complete: the function that can be approximated by the network is fixed, as the parameters in the corresponding UAT are fixed. So to ensure the network can dynamically approximate functions based on input, some or all of the parameters $\alpha_{j}$, $W_{j}$ and $\theta_{j}$ must be influenced by the input (referencing to UAT2LLMs and UAT2CVs), which can be simply written as $\alpha_{j}(\mathbf{x})$, $W_{j}(\mathbf{x})$ and $\theta_{j}(\mathbf{x})$. This allows the network to dynamically adapt and fit the function based on the input while enabling parallel computation. So we could design the basic block in Figure 1, like $\text{Block}_{1}\cdots\text{Block}_{n}$ are the implementation of $\alpha_{j}(\mathbf{x})\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}(\mathbf{x})\mathbf{x}+\theta_{j}(\mathbf{x})\right),i=1\cdots n$. Some or all the parameters $\alpha_{j}$, $W_{j}$ and $\theta_{j}$ will change according to input.
@@ 3.1 Requirements for Basic Block in Parallel Network
Figure 3: The structure of the Para-Former-1-n represents Para-Former with 1 depth of n layers.
Figure 4: The structure of the Para-Former-m-n represents Para-Former with m depth of n layers.
Cifar-10 Cifar-100 STL-10 OxfordIIITPet Flower-102
@@
3.2 Para-former
$=$ $\operatorname{Concat}\left(\hat{\mathbf{H}}_{i1},\ldots,\hat{\mathbf{H}}_{ih}\right)\mathbf{W}_{O}$
$\hat{\mathbf{H}}_{ij}$ $=\operatorname{Attention}(\mathbf{x}_{ij}\mathbf{W}_{Qij}=\mathbf{Q}_{ij},\mathbf{x}_{ij}\mathbf{W}_{Kij}=\mathbf{K}_{ij},$ (4)
$\mathbf{x}_{i}\mathbf{W}_{Vi}=\mathbf{V}_{ij})$
$=\operatorname{softmax}\left(\frac{\mathbf{Q}_{ij}\mathbf{K}_{ij}^{T}}{\sqrt{M}}\right)\mathbf{V}_{ij}$
$=\mathbf{H}_{ij}\mathbf{V}_{ij}=\mathbf{H}_{ij}[\mathbf{x}_{ij}\mathbf{W}_{Vij}]$
where the subscript of $ij$ represents the $j$-th head of Multi-Head Attention (MHA) in the $i$-th layer. Please note that throughout the following text, all the subscript $i$ refer to the layer index in the network, like $\mathbf{x}_{i}$. According to Figure 2, assuming the input to the Feed-Forward Network (FFN) in the $i$-th layer of the Transformer, it can be expressed as:
The above representation mainly describes the engineering process rather than its true mathematical form. According to UAT2LLMs, the MHA and FFN components of Transformers can be expressed in matrix-vector form. We denote the matrix-vector form of variables and operations with a prime symbol (e.g., $\mathbf{x}_{i}$ becomes $\mathbf{x}^{\prime}_{i}$). So MHA and FFN can be written as (For more details, refer to UAT2LLMs):
$MHA^{\prime}(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}$ (6)
$FFN^{\prime}(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{b}^{\prime}_{i+1,3}\$ (7)
According to Eq. (6) and Eq. (7), the output $\mathbf{x}^{\prime}_{1}$ of a single-layer Transformer can be expressed as Eq. (8) (see Appendix A for more details). The mathematical form of Eq. (8) can be considered equivalent to the two terms $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$ in Eq. (1). Therefore, if we parallelize multiple instances of $\mathbf{x}^{\prime}_{1}$, we obtain a multi-layer parallel Transformer network, as illustrated in Figure. 3.
$\mathbf{x}^{\prime}_{1}$ $=(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,2})$ (8)
Additionally, we provide the output for a two-layer Transformer, as shown in Eq. (9) (see Appendix A for more details). From Eq. (9), it is evident that the serial configuration of two Transformer layers also adheres to the mathematical form of UAT. Therefore, we can extend the Para-Former structure. Specifically, a block in Figure. 1 can contain multiple serial Transformer modules (it has been proven in UAT2LLMs that the mathematical form of multi-layer Transformers is also consistent with UAT), while multiple blocks operate in parallel, as illustrated in Figure. 4. The number of Transformer modules within a single block is referred to as the depth, while the total number of Transformer modules across all blocks is referred to as the network’s layer count.
$\mathbf{x}^{\prime}_{2}$ $=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (9)
$+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,2})$
Both design strategies possess the capability to dynamically adjust model parameters based on the input. In Figure 3, the primary mechanism is the dynamic adjustment of MHA weight parameters based on the input. In contrast, Figure 4 not only leverages the dynamic adjustment of MHA weight parameters but also utilizes UAT to dynamically approximate some bias parameters. For instance, in Eq. (9), $\mathbf{b}^{\prime}_{2,2}=(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$, can be seen as a two-layer UAT. Refer to Appendix B to see which parameters are influenced by the input. So, what are advantages of these strategies? We will talk about these in Section 4.3.
Para-F-1-1 Para-F-1-6 Para-F-1-12 Para-F-1-24